ELECTRONIC SPECTROSCOPY OF THE A^{1}A′′ − X^{1}A′ SYSTEM OF DCF

H. FAN, I. IONESCU, \underline{C. TAO}, M. DESELNICU, C. MUKARAKATE, AND S. A. REID, Department of Chemistry, Marquette University, Milwaukee, WI 53233; JU XIN, Department of Physics and Engineering Technologies, Bloomsburg University, Bloomsburg, PA 17815.

To further investigate the Renner-Teller effect and excited state barriers to linearity and dissociation in the simplest singlet carbene, HCF, we measured fluorescence excitation spectra and lifetimes of the pure bending levels 2_{0}^{n} and combination bands 1_{0}^{1}2_{0}^{n} and 2_{0}^{n}3_{0}^{1} in the A^{1}A′′-X^{1}A′ system of the deuterated isotopomer, DCF. The spectra were measured under jet-cooled conditions using a pulsed discharge source, and rotationally analyzed to yield precise values for the band origins and rotational constants. The derived A^{1}A′′ state parameters are compared with our previous results for HCF, and with the predictions of ab initio electronic structure theory.